OpenStudy (anonymous):

MEDALS FOR WHO SOLVES THE EQUATIONS! Graphing or substituting for a linear systems of equations. y=5-3x x-y=7 Answer: (3,-4) How do you solve to get the answer. Show work

OpenStudy (nikato):

use substitutions. plug the first equation into teh second such that you have x-(5-3x)=7 then solve for x

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

x-(5-3x)=7 then solve for x i was going to say that

OpenStudy (nikato):

you mean *x=3 now plug this value of x into any one of the 2 equations to solve for y

OpenStudy (jdoe0001):

\(\bf {\color{red}{ y}}=5-3x\\ \quad \\ x-{\color{red}{ y}}=7 \implies x-({\color{red}{ 5-3x}})=7\) <--- substitution, thus

OpenStudy (anonymous):

Yes I do mean x.

OpenStudy (anonymous):

can you gimme one medal too

OpenStudy (nikato):

so... you can either plug into the first equation y=5-3(3) OR the second 3-y=7 and solve for y

OpenStudy (anonymous):

y=-4

OpenStudy (anonymous):

you got it

OpenStudy (nikato):

okey dokey. done..

OpenStudy (anonymous):

thanks cassieblake

OpenStudy (anonymous):

:) no problem. Now it is far.